# Properties of Carbon-rich AGB Stars in the LMC and the Milky Way

### v2

@@ 4.2 SEDs of Galactic CAGB stars
4.3 Distances of Galactic CAGB stars
For Galactic CAGB stars, the top panels of Figure 7 compare distances derived from Gaia DR3 parallax measurements with those obtained from theoretical CAGB model SEDs, shown separately for MW-CAGB_IC (left) and MW-CAGB_NI (right) stars. Although considerable scatter is present, the two distance estimates are generally consistent, with MW-CAGB_IC objects showing better overall agreement. The upper-left panel also highlights a comparison for the MW-CAGB_IC-S subgroup, a subset of CAGB_IC stars (see Section 4.2).
Of the 4909 MW-CAGB_IC objects (see the upper-left panel of Figure 7), 78.6% fall within a $\pm$0.23 dex agreement between the two distance estimates, as indicated by the lime dotted lines in the panel. This corresponds to distances agreeing within a factor of 1.7 (approximately -41% to +70% difference). When restricting the sample to the 114 objects with distances from the CAGB model SEDs within 900 pc, the agreement improves to 84.2%. Of the 2254 MW-CAGB_NI objects (see the upper-right panel of Figure 7), 55.9% fall within a $\pm$0.23 dex agreement between the two distance estimates.
The middle panels of Figure 7 compare distances derived from the period–magnitude relation (PMR) calibrated with LMC Miras (see Section 6; Equation 1) with those obtained from CAGB model SEDs for Mira-type CAGB stars. The two methods show strong consistency: among 1035 Mira-type MW-CAGB_IC stars (middle-left panel), 84.4 % agree within $\pm$0.23 dex, while for 114 Mira-type MW-CAGB_NI stars (middle-right panel), 71.1 % fall within the same range. Although both approaches have inherent limitations and uncertainties, SED fitting and the PMR provide the only practical means of deriving reliable distances from observed photometric and variability data for large samples of CAGB stars.
The lower-left panel of Figure 7 shows the WISE–2MASS CMD for Galactic CAGB stars, using distances derived from Gaia DR3 parallaxes. The diagram exhibits considerable scatter, with many absolute magnitudes lying outside the luminosity-function range for CAGB stars discussed in Section 3.2, and falling within the region occupied exclusively by extrinsic carbon stars, as identified by Suh (2024).
@@ 4.3 Distances of Galactic CAGB stars
Figure 7: For Galactic CAGB stars, the upper 4 panels show comparisons between distances derived from the CAGB model SEDs, Gaia DR3 parallax measurements, and the PMR obtained from Miras in the LMC. The bottom panels show IR CMDs for Galactic CAGB stars using distances from Gaia DR3 (left) and CAGB model SEDs (right). For CAGB models (AMC, $T_{c}$ = 1000 K): $\tau_{10}$=0.0001, 0.01, 0.1, 0.5, 1, 2, and 4 from left to right (see Table 3). The error bars represent the uncertainties in the derived distances. The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 4.3 for details.
Figure 8: The top and lower-left panels show IR 2CDs for CAGB stars in the LMC and the Milky Way, while the lower-right panel compares fluxes in the W4[22] and S5[24] bands. For CAGB models (AMC, $T_{c}$ = 1000 K): $\tau_{10}$=0.0001, 0.01, 0.1, 0.5, 1, 2, and 4 from left to right (see Table 3). The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 5.
Figure 9: IR CMDs for CAGB stars in the LMC and the Milky Way. For CAGB models (AMC, $T_{c}$ = 1000 K): $\tau_{10}$=0.0001, 0.01, 0.1, 0.5, 1, 2, and 4 from left to right (see Table 3). For Galactic CAGB stars, distances are derived from CAGB model SEDs with Gaussian Monte Carlo scatter (GS) in $\log_{10}d$ ($\sigma$$\pm$0.0798 dex; obtained from LMC-CAGB stars; see Figure 1). The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 5.
5 2CDs and CMDs in IR Bands
@@
A comparable patten appeared in the WISE–2MASS CMDs (Figures 2 and 3), which compares CAGB stars in the LMC and the Milky Way, plotting M(W3[12]) and M(K[2.2]) against K[2.2]-W3[12]. Likewise, Figure 9 presents Spitzer–WISE–2MASS CMDs extending to longer wavelengths (W4[22] and S5[24]). The left panel shows M(S5[24]) versus K[2.2]-W3[12] for LMC-CAGB stars, while the right panel displays M(W4[22]) versus K[2.2]-W3[12] for Galactic CAGB stars.
@@
6 IR properties of Mira variables
@@
Figure 10: PMRs for Mira-type CAGB stars. The top panel shows the PMR for LMC-CAGB objects and the bottom panel shows the PMR for MW-CAGB objects. The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 6.
Figure 11: The upper two panels compare distances derived from the LMC-based PMR with those from Gaia DR3 for Mira-type Galactic CAGB stars. The bottom panel compares distances derived from the LMC-based PMR with those obtained from CAGB model SEDs for Mira-type LMC-CAGB stars. The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 6.

### v4

@@ IV.2 SEDs of Galactic CAGB stars
IV.3 Distances of Galactic CAGB stars
For Galactic CAGB stars, the top panels of Figure 7 compare distances derived from Gaia DR3 parallax measurements with those obtained from theoretical CAGB model SEDs, shown separately for MW-CAGB_IC (left) and MW-CAGB_NI (right) stars. Although considerable scatter is present, the two distance estimates are generally consistent, with MW-CAGB_IC objects showing better overall agreement. The upper-left panel also highlights a comparison for the MW-CAGB_IC-S subgroup, a subset of CAGB_IC stars (see Section IV.2).
Of the 4909 MW-CAGB_IC objects (see the upper-left panel of Figure 7), 78.6% fall within a $\pm$0.23 dex agreement between the two distance estimates, as indicated by the lime dashed lines in the panel. This corresponds to distances agreeing within a factor of 1.7 (approximately -41% to +70% difference). When restricting the sample to the 114 objects with distances from the CAGB model SEDs within 900 pc, the agreement improves to 84.2%. Of the 2254 MW-CAGB_NI objects (see the upper-right panel of Figure 7), 55.9% fall within a $\pm$0.23 dex agreement between the two distance estimates.
The middle panels of Figure 7 compare distances derived from the period–magnitude relation (PMR) calibrated with LMC Miras (see Section VI; Equation 1) with those obtained from CAGB model SEDs for Mira-type CAGB stars. The two methods show strong consistency: among 1035 Mira-type MW-CAGB_IC stars (middle-left panel), 84.4 % agree within $\pm$0.23 dex, while for 114 Mira-type MW-CAGB_NI stars (middle-right panel), 71.1 % fall within the same range. Although both approaches have inherent limitations and uncertainties, SED fitting and the PMR provide the only practical means of deriving reliable distances from observed photometric and variability data for large samples of CAGB stars.
The lower-left panel of Figure 7 shows the WISE–2MASS CMD for Galactic CAGB stars, using distances derived from Gaia DR3 parallaxes. The diagram exhibits considerable scatter, with many absolute magnitudes lying outside the luminosity-function range for CAGB stars discussed in Section III.2, and falling within the region occupied exclusively by extrinsic carbon stars, as identified by Suh (2024).
@@ IV.3 Distances of Galactic CAGB stars
Figure 7: For Galactic CAGB stars, the upper 4 panels show comparisons between distances derived from the CAGB model SEDs, Gaia DR3 parallax measurements, and the PMR obtained from Miras in the LMC. The bottom panels show IR CMDs for Galactic CAGB stars using distances from Gaia DR3 (left) and CAGB model SEDs (right). For CAGB models (AMC, $T_{c}$ = 1000 K): $\tau_{10}$=0.0001, 0.01, 0.1, 0.5, 1, 2, and 4 from left to right (see Table 3). The error bars represent the uncertainties in the derived distances. The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section IV.3 for details.
Figure 8: The top and lower-left panels show IR 2CDs for CAGB stars in the LMC and the Milky Way, while the lower-right panel compares fluxes in the W4[22] and S5[24] bands. For CAGB models (AMC, $T_{c}$ = 1000 K): $\tau_{10}$=0.0001, 0.01, 0.1, 0.5, 1, 2, and 4 from left to right (see Table 3). The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section V.
Figure 9: IR CMDs for CAGB stars in the LMC and the Milky Way. For CAGB models (AMC, $T_{c}$ = 1000 K): $\tau_{10}$=0.0001, 0.01, 0.1, 0.5, 1, 2, and 4 from left to right (see Table 3). For Galactic CAGB stars, distances are derived from CAGB model SEDs with Gaussian Monte Carlo scatter (GS) in $\log_{10}d$ ($\sigma$$\pm$0.0798 dex; obtained from LMC-CAGB stars; see Figure 1). The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section V.
V 2CDs and CMDs in IR Bands
@@
A comparable patten appeared in the WISE–2MASS CMDs (Figures 2 and 3), which compare CAGB stars in the LMC and the Milky Way, plotting M(W3[12]) and M(K[2.2]) against K[2.2]-W3[12]. Likewise, Figure 9 presents Spitzer–WISE–2MASS CMDs extending to longer wavelengths (W4[22] and S5[24]). The left panel shows M(S5[24]) versus K[2.2]-W3[12] for LMC-CAGB stars, while the right panel displays M(W4[22]) versus K[2.2]-W3[12] for Galactic CAGB stars.
@@
VI IR properties of Mira variables
@@
Figure 10: PMRs for Mira-type CAGB stars. The top panel shows the PMR for LMC-CAGB objects and the bottom panel shows the PMR for MW-CAGB objects. The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section VI.
Figure 11: The upper two panels compare distances derived from the LMC-based PMR with those from Gaia DR3 for Mira-type Galactic CAGB stars. The bottom panel compares distances derived from the LMC-based PMR with those obtained from CAGB model SEDs for Mira-type LMC-CAGB stars. The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section VI.